40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2253     Entry time: Sun Nov 18 18:50:39 2018
Author: anchal 
Type: Summary 
Category: South Cavity 
Subject: South Path up and running again 

The South PMC is installed and the South FSS is running again. Attached is the beam path profile from ala mode used for the south path.
While aligning this beam into the cavity (which is a very difficult task apparently), I disturbed the north path by mistake. So I had to align both paths from scratch again.

Few notes for aligning beam into cavities in future:

  • The only way to know if the beam is reflecting off the mirror, not the enclosure or something else inside the tank, is to bend and look through an IR viewer. So my first step would be to do this and ensure the beam is at least hitting the mirror. This step requires the removal of the thermal shield.
  • After this, it takes a while to actually start seeing the reflection from the mirror. I bend to see into the cavity to see if the reflection from the mirror is hitting something else and change angles accordingly.
  • At this stage, once we get the reflection, we very nicely just match it with the incident beam. BUT, make sure you are not seeing anything in the transmission. It is possible that the beam is simply going through the spacer. A way to check if the transmitted light is through the cavity or spacer is by changing laser frequency (Slow input to laser) slightly.
  • Once we get the coincident reflection with no transmission, scan the slow input about 0.75 V to see if any of the modes are excited. If you get lucky, some lower order mode gets excited to give some information about how wrong the incident beam is.
  • From this point onwards, it is just walking of beam all over the mirror of the cavity and scanning laser frequency to get any glimpse of the fundamental mode.
  • Once the fundamental mode is caught, we actuate the fast input to the laser with ~3 Vpp sinusoidal signal at around 4 Hz. Adjust the periscope to get some tangible transmission and then just do the usual optimization by walking the beam.

Important things to remember:

  • Make the autolock condition of PMC less hard so that it remains locked with the fluctuating laser frequency.
  • Every time you are moving a knob, double check if you are moving the right one. This is how I fucked up a perfectly good alignment on the North side.

Also, I have made correct length cables for FSS LO delays for both paths and both cavities are getting locked nicely. At North cavity, I see ~70% mode matching and at South cavity ~60%. These were measured with laser power meter.

Future steps are to align ouput of the cavities onto Beat note detector and start measuring noisebudget weekly. After that, we'll concentrate on increasing mode matching and reducing noise sources in the paths.

Attachment 1: SouthOptimizedPath.pdf  105 kB  | Hide | Hide all
SouthOptimizedPath.pdf
ELOG V3.1.3-